Sam found a number of nickels, dimes, and quarters inhis room. He found 6 more dimes than mickels but three times as many quarters as dimes. The total value of the coins was $11.40. howmany coins of each type did Sam find?
Answer by Alan3354(69443)

Sam found a number of nickels, dimes, and quarters inhis room. He found 6 more dimes than mickels but three times as many quarters as dimes. The total value of the coins was $11.40. howmany coins of each type did Sam find?
--------------
d = n+6
q = 3d --> q = 3n+18
25q + 10d + 5n = 1140
25(3n+18) + 10(n+6) + 5n = 1140
75n+450 + 10n+60 + 5n = 1140
90n + 510 = 1140
90n = 630
n = 7
d = 13
q = 39